# The unrealized potential of agroforestry for an emissions-intensive agricultural commodity

### v1

@@ Methods
We will describe our methods for a threshold level of cover of 30 %, noting that we used the same procedure for all threshold levels. For any one hypothetical pixel, we first use our fitted regression model to estimate the biomass associated with a cover level of 30%. This was done by randomly drawing 100 samples of paired intercept and slope estimates from the posterior distributions of our fitted regression model, and then using each of these unique parameter combinations to generate a posterior distribution of biomass for a hypothetical pixel with 30% cover. We then multiplied this posterior distribution by the total number of pixels in our cover map that have less than 30% shade-tree cover, which gives a posterior distribution of the biomass that would be added by increasing the shade cover of those pixels to 30%. We stress that this method replaces our estimate of the biomass within each pixel below the threshold value (from our AGBD map) with the amount of biomass that would be present if cover was increased to the threshold value, while accounting for uncertainty in the underlying relationship between cover and biomass. We then added the total biomass of all pixels already at or above the threshold level by extracting these values directly from our biomass map. Our final biomass estimate is a posterior distribution of total biomass across all locations where shade tree cover is increased to 30% wherever it is below that threshold, and remains unchanged for all locations where the threshold is currently satisfied. From this distribution, we extracted both the mean total carbon and a 95% Highest Density Interval (HDI), representing the range within which the mean biomass is statistically most likely to reside. Biomass was converted to carbon assuming that carbon represents 47% of the above-ground biomass, following Martin et al. (2011) [55]. The same procedure was repeated for all threshold levels.
@@
Data availability
The maps of shade-tree cover and map of aboveground biomass density will be released for download and will be available in the Google Earth Engine. The ground truth data will be made available in an online data repository.
Code availability
The code to estimate shade tree cover and above-ground biomass density will be made available on github.
Acknowledgements
We thank R. Tetteh, J. Afele, E. Nimo, M. Yombu, V. Agumenu, and Hammond for their exceptional efforts in visiting hundreds of cocoa farms to collect the ground truth data in Ghana. Our thanks also extend to Barry Callebaut and Olam Food Ingredients (ofi) for their invaluable support in making the ground campaign possible. Special thanks to M. Gilmour, S. Ankamah, C. Parra Paitan, O. Nkuah, E. Prempeh, R. Seidu, B. Karibu, S. Adusei, S. Dodzie, E. Obiri Yeboah, S. Larbie, and D. Forson and their ground-level field teams for their support in helping us gain access to recently mapped cocoa farms . We are grateful to the German Development Agency (GIZ) for their collaboration, which enabled us to extend our ground truth sampling to Côte D’Ivoire. Special thanks to H. Walz, A. Bio, P. Ripplinger, and M. Pallauf. We also appreciate the support of P. Kouakou in processing high-resolution drone images for Côte d’Ivoire. Thanks to Dr. Alexi Ernstoff, Vincent Rossi, Cecile Guignard, and Tereza Levova from Quantis for their assistance with analysis of annual carbon emissions from cocoa production. This project received funding from the Lindt Cocoa Foundation (W.J.B., S.P.H, J.D.W), the 2019–2020 BiodivERsA joint call for research proposals under the BiodivClim ERA-Net COFUND program, with the funding organization of the Swiss National Science Foundation (R.D.G., C.B., J.D.W. and W.J.B.), the Joint Cocoa Research Fund of CAOBISCO and ECA (W.J.B., S.P.H, J.D.W), and the Queensland Government under the Women’s Research Assistance Program (WRAP194-2019RD1; W.J.B.). We greatly appreciate the open data policies of the ESA Copernicus program.
Author contributions statement
W.J.B., S.P.H., and J.D.W. conceived the research question and designed the study. W.J.B. planned the fieldwork, with W.J.B. and E.D. overseeing its execution. A.B. developed the code with guidance from J.D.W. and K.S. W.J.B., A.B., and S.P.H. analyzed the results. C.B. and F.C. created the maps of agro-climatic zones. W.J.B. and S.P.H. led the writing and wrote the first draft of the manuscript with inputs from A.B and J.D.W., and all authors contributed to revisions.
Supplementary methods
Comparison to existing AGBD estimates
We compare our biomass map with the European Space Agency’s (ESA) Climate Change Initiative (CCI) state-of-the-art Biomass map[31] for 2018. The ESA map is intended to provide global coverage of AGBD estimates with a ground sampling distance of $100\times 100$ meters (i.e., four times less pixels per area than our map), with the aim of monitoring global biomass change in the context of climate change. The ESA map is created bi-annually based on multi-modal input data like spaceborne SAR (newer versions also include GEDI) and processed using a proprietary algorithm. Because we are lacking an independent, large-scale ground truth dataset of AGBD collected in the field across our region of interest, we compare both maps to 44,200 GEDI footprints that fall in our hold-out test areas (i.e., unseen by our model during training and validation) and contain cocoa according to Kalischek et al.[45]. One should keep in mind though that our map has been specifically fine-tuned to our area of interest (as opposed to the ESA CCI map, which is a global product) and it is using GEDI footprint-level AGBD data for training, too. Unknown global biases that could be present in the GEDI AGBD footprint-level product would thus most likely impact our map but not show up in the comparison to the hold-out GEDI footprints here, if those would exhibit the same kind of possible bias. This comparison shall therefore not judge the quality of the ESA CCI map but rather provide some rough notion about our map’s properties. Overall, our map achieves (all values in tonnes ha-1) a mean absolute error (MAE) of 36.2 (vs. ESA’s 56.9), a root mean squared error (RMSE) of 63.2 (vs. ESA’s 87.9) and a mean error (bias) of 6.5 (vs. ESA’s 33.11). This shows that, in cocoa areas, our map does not only exhibit lower errors than the comparison map, but is also less biased towards lower AGBD values. In Fig. S5b, we additionally show an analysis of the residuals between the two maps and the GEDI reference footprints, to better understand the biases of both maps over different AGBD intervals.
Description of agro-climatic zones in Ghana and Côte d’Ivoire.
Agro-climatic zoning (AEZ) is a common tool to identify analogous environments to prioritize interventions. Cocoa is limited to the tropical wet & dry (Aw) and tropical wet (Ar) zones but does not fully occupy these two climate zones [34]. We therefore developed agro-climatic zones by identifying groups with common climate characteristics in known cocoa growing areas using the RandomForests (RF) [20] classifier similar to Bunn et al. [21]. However, in contrast to Bunn et al. [21] we sampled cocoa occurrences from Kalischek et al. [45] and clustered them on climate data to identify differnt agro-climatic types. The RF classifier was then trained and extrapolated on climate data to generate maps of the agro-climatic zones.
Occurrence data. We obtained cocoa occurrence from high-resolution maps of cocoa plantations for Côte d’Ivoire and Ghana [45]. This data consisted of cocoa plots identified on Copernicus Sentinel-2 optical satellite imagery using a deep learning framework and <100.000 GPS polygon mapped cocoa farms. We first geographically subsampled at 2.5’ from this layer. We obtained <9000 unique cocoa locations, from which we drew a 10% random sample as reference distribution.
Climate analysis. We constructed a monthly climatology from climatic data from CHIRTS ERA-5 [38] and CHIRPS v2.0 [39], encompassing precipitation and temperature data from January 1990 to December 2020. These data were originally available at a daily temporal resolution and 0.05° spatial resolution. Additionally, we incorporated solar radiation data obtained from the Global PET and Aridity Index database [82]. From the monthly climatology, we derived biologically 33 meaningful variables (Table S1). To the commonly used set of 19 bioclimatic variables, we added variables that directly relate to cocoa cultivation. As as described in Bunn et al.[21], a short dry season may reduce overall yields, excessive growing season temperatures may affect quality, and low vapour pressure may reduce assimilation rates. We therefore added the number of consecutive months with less than 100mm of precipitation, the number of consecutive arid months (precipitation lower than potential evapotranspiration (PET)), the total water deficit during this period, growing season average temperatures and mean dry season maximum temperatures as additional bioclimatic variables. Potential evapotranspiration is considered a good proxy of vapour pressure [2] which may have direct effects on the cocoa tree [56]. Based on precipitation, solar radiation, maximum temperature, mean temperature, and minimum temperature data, we estimated potential evapotranspiration (ETps) following the modified Hargreaves approach as recommended by the FAO [2]:
$\text{ETps}=0.0013\times 0.408\times\text{Srad}\times(\text{Tavg}+17)\times\left(\text{Tmax}-\text{Tmin}-0.0123\times\text{Prec}\right)^{0.76}$ (1)
where Srad represents solar radiation, Tavg is the average temperature, Tmax and Tmin denote the maximum and minimum temperatures, respectively, and Prec stands for precipitation. We initially considered all 33 bioclimatic variables (Table S1). However, to reduce multicollinearity and enhance model reliability, we applied a Variance Inflation Factor (VIF) analysis. This process identified a subset of 13 variables that were used in the final modeling to define agro-climatic types (Table S1).
Table S1: Available bioclimatic variables and variables used to model the distribution of suitable climates for cocoa production across Ghana and Côte d’Ivoire.
Type Variable Description Included in model
Temperature BIO 1 Annual mean temperature
BIO 2 Mean diurnal range (mean of monthly (max temp - min temp))
BIO 3 Isothermality (BIO2/BIO7) (*100) x
BIO 4 Temperature seasonality (standard deviation *100) x
BIO 5 Max temperature of warmest month
BIO 6 Min temperature of coldest month
BIO 7 Temperature annual range (BIO5-BIO6)
BIO 8 Mean temperature of wettest quarter x
BIO 9 Mean temperature of driest quarter
BIO 10 Mean temperature of warmest quarter
BIO 11 Mean temperature of coldest quarter x
Precipitation BIO 12 Annual precipitation
BIO 13 Precipitation of wettest month
BIO 14 Precipitation of driest month x
BIO 15 Precipitation seasonality (coefficient of variation) x
BIO 16 Precipitation of wettest quarter
BIO 17 Precipitation of driest quarter
BIO 18 Precipitation of warmest quarter x
BIO 19 Precipitation of coldest quarter x
BIO 20 Number of consecutive months < 100mm precipitation x
Growing Season BIO 21 Consecutive months with less prec than PET
BIO 22 Sum of water deficit during dry season x
BIO 23 Mean temperature during wettest semester
BIO 24 Max dry season temperature
Evapotranspiration BIO 25 Annual PET x
BIO 26 PET seasonality (coefficient of variation)
BIO 27 Max PET
BIO 28 Min PET
BIO 29 Range of PET (Max PET - Min PET)
BIO 30 PET of wettest quarter x
BIO 31 PET of driest quarter
BIO 32 PET of warmest quarter x
BIO 33 PET of coldest quarter
Definition of agro-climate types. We used the Random Forest (RF) [20] classifier in two distinct applications. (1) We initially used it to produce a dissimilarity measure to group occurrence locations into agro-climatic clusters with similar climate characteristics in an unsupervised variation. (2) We used the RF classifier to classify climate data of current and future conditions into the resulting agro-climatic types. Random Forests are machine learning classifiers that are formed by ensembles of classification trees. RF are very popular because of their efficiency on large datasets without overfitting. We used the randomForest package [52] in the statistical software R [64] that implements the RF approach. First, to determine distinct agro-climatic types for cocoa we used the RF classifier in unsupervised mode [75] to calculate dissimilarities based on the bioclimatic variables at cocoa locations. The RF measure of dissimilarity that results in biologically meaningful groupings when used as input for clustering instead of common Euclidean distances [75]. To increase cluster stability, we used features with low pair-wise correlation, a high number of forests (100) and trees (1000), and specified variables picked at each node (3). Clustering was performed using Ward hierarchical clustering using the dissimilarities as input. The number of clusters was determined based on visual inspection of a cluster dendrogram. To train the RF classifier to recognize the resulting suitable climate types, we extracted the values of the 13 selected climate variables using the geolocation data of the cocoa occurrences. In addition, a random background sample set of points not known to produce cocoa was drawn to characterize the general environment at a sampling ration of 1:1 background to occurrence locations. For binary classification problems such a sampling ratio is recommended to avoid the preferential prediction of the majority class [11]. The sampling ratio we chose accounted for the trade-off between the multi-class suitability classification, as well as the binary classification in to suitable classes and unsuitable background locations. Additionally, we constrained the background samples to contain only pixels within cocoa producing countries to exclude unfeasible locations [85]. Each RF forest classifier was configured to create 100 decision trees, and to be replicated 25 times. We followed the recommendation to increase the number of trees until the prediction error of independent runs stabilizes [18]. In each repeat, a different subset from the occurrence sample was drawn at the size equaling half the number of cases in the smallest subgroup. Such reduction of ecological sampling bias has been shown to improve the capacity of niche-based approaches to correctly predict species distributions [86]. From this subset, we used 80% for training, and 20% for evaluation. With this configuration and inputs, the classifier was extrapolated onto maps of the 13 bioclimatic variables and determined the modal value across all trees and forests. The modal classification of all RF trees was used to determine the corresponding zonal distribution of agro-climatic types under current climatic conditions using all climate variables. In addition, we used the percentage of votes by individual trees in the forests to evaluate the model agreement. Locations that received similar votes for multiple suitable classes were labeled as "mixed" to reflect the ambiguity in classification. For locations where the majority of votes indicated "unsuitable," but a significant number of votes for suitable classes were still present, these were labeled as "Limitations."
Results. We identified five distinct agro-climatic types across Ghana and Côte d’Ivoire, each showing unique patterns of cocoa suitability. For each climate variable, we compared the confidence intervals for each climate type resulting from RF classification relative to all other clusters (Figure S2). Notable differences of each type relative to the other occurrence locations were used to describe the climate types as follows:
  1. 1.
    Type 1 - Moderate – very dry: Intermediate values for most temperature variables (Bio 3, 4, 8), but low values for precipitation during the driest month (Bio 14) and high number of consecutive months with less than 100 mm of precipitation (Bio 20).
  2. 2.
    Type 2 - Hot – dry: Relatively high temperature values (Bio 8 and 11), high precipitation seasonality (Bio 15), long dry season (Bio 20).
  3. 3.
    Type 3 - Hot – wet: Relatively high temperature values (Bio 8 and 11), high precipitation during the driest month (Bio 14), lowest number of consecutive dry months (Bio 20).
  4. 4.
    Type 4 - Cold – wet: Low mean temperature of the wettest quarter (Bio 8), high precipitation of the coldest quarter (Bio 19), relatively high number of consecutive dry months (Bio 20), high potential evapotranspiration (Bio 25).
  5. 5.
    Type 5 - Very hot – very wet: High temperature seasonality and values (Bio 4, 8, and 11), low number of consecutive dry months (Bio 20).
These five agro-climatic types were mapped as five agro-climatic zones for Ivory Coast and Ghana, along with three additional classifications: "unsuitable" zones, which are unlikely to support cocoa due to climatic constraints, "mixed" zones which could not be clearly categorized into one of the agro-climatic zones and "limitations" zones which were classified unsuitable but were found to contain cocoa occurrences (Figure S1). This mapping provides a comprehensive view of the suitability of different areas for cocoa cultivation, considering the varying constellations of climatic attributes. Briefly, the Type 1 – moderate- very dry zone spanned the Northern cocoa zone from Yamoussoukro (CdI) to Ahafo (Ghana). Type 2 – hot -dry climate can be found South of Type 1 and North of Type 3 from Goh (CdI) to Ashanti (Ghana). The central cocoa zones Soubre and Divo in CdI, as well as Western Region in Ghana were characterized by Type 3 - hot- wet climate. Type 4 – cold – wet – climate can be found in the Montagnes area (CdI) and the Ghanaian boarder with Togo, which also has some higher elevation area. Type 5 – very hot – very wet climate mostly covered coastal areas along the Gulf of Guinea (Figure S1).

### v2

@@ Methods
We will describe our methods for a threshold level of cover of 30%, noting that we used the same procedure for all threshold levels. For any one hypothetical pixel, we first use our fitted regression model to estimate the biomass associated with a cover level of 30%. This was done by randomly drawing 100 samples of paired intercept and slope estimates from the posterior distributions of our fitted regression model, and then using each of these unique parameter combinations to generate a posterior distribution of biomass for a hypothetical pixel with 30% cover. We then multiplied this posterior distribution by the total number of pixels in our cover map that have less than 30% shade-tree cover, which gives a posterior distribution of the biomass that would be added by increasing the shade cover of those pixels to 30%. We then added the total biomass of all pixels already at or above the threshold level by extracting these values directly from our biomass map. Our final biomass estimate is a posterior distribution of total biomass across all locations where shade tree cover is increased to 30% wherever it is below that threshold, and remains unchanged for all locations where the threshold is currently satisfied. From this distribution, we extracted both the mean total carbon and a 95% Highest Density Interval (HDI), representing the range within which the mean biomass is statistically most likely to reside. Biomass was converted to carbon assuming that carbon represents 47% of the above-ground biomass, following Martin et al. (2011) [martin2011]. The same procedure was repeated for all threshold levels. Additional details of our carbon sequestration calculations are described in the Supplementary Methods.
Data availability
All key datasets used in this study are publicly available. The final maps of shade-tree cover, canopy height, and aboveground biomass density can be explored interactively and downloaded via Google Earth Engine: https://albecker.users.earthengine.app/view/agroforestry. The drone-derived ground-truth data for shade-tree cover are available via UQ eSpace https://doi.org/10.48610/dda018c[blaser-hart2025cocoa]. Sentinel-2 satellite imagery is available from the Copernicus Open Access Hub: https://sentinels.copernicus.eu. Aboveground biomass reference data were obtained from NASA GEDI L4A (2022), available at https://doi.org/10.3334/ORNLDAAC/2056. Canopy height data were generated using the model of Lang et al. (2023)[lang2023high]. The cocoa-growing area mask used in this study was obtained from Kalischek et al. (2023)[kalischek23natfood] and used as provided. Forest classifications were based on the JRC Tropical Moist Forest (TMF) map: https://forobs.jrc.ec.europa.eu/TMF/. Climatic variables were derived from the CHIRTS-daily temperature dataset and the CHIRPS v2.0 precipitation dataset: https://www.chc.ucsb.edu/data. Potential evapotranspiration (PET) data were obtained from the Global Aridity and PET Database: https://cgiarcsi.community/data/global-aridity-and-pet-database/. Administrative boundaries were sourced from GADM v4.1: https://gadm.org. The greenhouse gas emission estimates for cocoa production were provided by Quantis using the World Food LCA Database, which is not publicly available; however, the exact values used in this study are reported in the manuscript. Farm boundary polygons used to guide drone flight planning were provided by cocoa trading companies and are not publicly available due to confidentiality agreements.
Code availability
The code used to estimate shade-tree cover and above-ground biomass density is available on GitHub: https://github.com/prs-eth/agroforestry. This repository includes code for pre-processing satellite imagery, training the gradient boosting and deep learning models (implemented in scikit-learn and PyTorch), and generating the final maps. The code used to calculate the climate-change mitigation scenarios is available via UQ eSpace https://doi.org/10.48610/dda018c[blaser-hart2025cocoa].
Corresponding author
Correspondence to Wilma J. Blaser-Hart.
Acknowledgements
We thank R. Tetteh, J. Afele, E. Nimo, M. Yombu, V. Agumenu, and Hammond for their exceptional efforts in visiting hundreds of cocoa farms to collect the ground truth data in Ghana. Our thanks also extend to Barry Callebaut and Olam Food Ingredients (ofi) for their invaluable support in making the ground campaign possible. Special thanks to M. Gilmour, S. Ankamah, C. Parra Paitan, O. Nkuah, E. Prempeh, R. Seidu, B. Karibu, S. Adusei, S. Dodzie, E. Obiri Yeboah, S. Larbie, and D. Forson and their ground-level field teams for their support in helping us gain access to recently mapped cocoa farms . We are grateful to the German Development Agency (GIZ) for their collaboration, which enabled us to extend our ground truth sampling to Côte D’Ivoire. Special thanks to H. Walz, A. Bio, P. Ripplinger, and M. Pallauf. We also appreciate the support of P. Kouakou in processing high-resolution drone images for Côte d’Ivoire. Thanks to Dr. Alexi Ernstoff, Vincent Rossi, Cecile Guignard, and Tereza Levova from Quantis for their assistance with analysis of annual carbon emissions from cocoa production. This project received funding from the Lindt Cocoa Foundation (W.J.B., S.P.H, J.D.W), the 2019–2020 BiodivERsA joint call for research proposals under the BiodivClim ERA-Net COFUND program, and with the funding organization of the Swiss National Science Foundation (FNRS under Grant n°PINT MULTI/BEJ—R.8002.20; R.D.G., C.B., J.D.W. and W.J.B.), the Joint Cocoa Research Fund of CAOBISCO and ECA (W.J.B., S.P.H, J.D.W), and the Queensland Government under the Women’s Research Assistance Program (WRAP194-2019RD1; W.J.B.). We thank the ESA Copernicus programme for its commitment to open data access, which made this research possible.
Author contributions
W.J.B., S.P.H., and J.D.W. conceived the research question and designed the study. W.J.B. planned the fieldwork, with W.J.B. and E.D. overseeing its execution. A.B. developed the code with guidance from J.D.W. and K.S. W.J.B., A.B., and S.P.H. analyzed the results. C.B. and F.C. created the maps of agro-climatic zones. W.J.B. and S.P.H. led the writing, with contributions from all authors.
Competing interests
The authors declare no competing interests.
Supplementary information
Supplementary methods
Agro-climatic zones in Ghana and Côte d’Ivoire.
Agro-climatic zoning (ACZ) is widely used to delineate comparable environmental regions in order to guide targeted interventions. No existing agro-climatic classification consistently covered both Ghana and Côte d’Ivoire in a crop-relevant way, so we developed a new scheme tailored to cocoa production systems across the region. To delineate agro-climatic zones, we grouped cocoa-growing areas with similar climate conditions using the Random Forest (RF) classifier [breiman2001], following the approach of Kamath et al. [Kamath2024]. However, in contrast to Kamath et al. [Kamath2024], we sampled cocoa occurrences from Kalischek et al.[kalischek23natfood] and clustered them using more recent climate data to identify distinct agro-climatic types. The RF classifier was then trained and extrapolated to generate maps of the agro-climatic zones.
Occurrence data. We obtained cocoa occurrence data from high-resolution maps of cocoa-growing areas in Côte d’Ivoire and Ghana [kalischek23natfood]. This data consisted of cocoa plots identified on Copernicus Sentinel-2 optical satellite imagery using a deep learning framework and <100.000 GPS polygon mapped cocoa farms. We first geographically subsampled at 2.5’ from this layer. We obtained <9000 unique cocoa locations, from which we drew a 10% random sample as reference distribution.
Climate analysis. We constructed a monthly climatology using temperature data from CHIRTS-ERA5 [funk2019] and precipitation data from CHIRPS v2.0 [funk2014], spanning January 1990 to December 2020. Both datasets were originally available at daily resolution and 0.05° spatial resolution. Solar radiation inputs were obtained from the Global PET and Aridity Index database [trabucco2018] and used to calculate potential evapotranspiration (PET) via the modified Hargreaves method [Allan1998].
To characterize climatic suitability for cocoa, we followed the approach of Kamath et al. [Kamath2024], deriving 33 bioclimatic variables: the standard 19 variables from WorldClim, along with cocoa-relevant metrics such as the number of consecutive dry months, arid periods (where precipitation is less than PET), total dry-season water deficit, growing season temperatures, and dry-season maximum temperatures. To reduce multicollinearity, we applied variance inflation factor (VIF) analysis, resulting in a final subset of 13 variables used to model agro-climatic zones (see Supplementary Fig. S3 for the selected variables).
Definition of agro-climate types. Following the methods outlined in Kamath et al. [Kamath2024], we used the Random Forest (RF) classifier [breiman2001] in two main steps: (1) identifying agro-climatic clusters in cocoa-growing regions, and (2), classifying current climate conditions into these types. RF was chosen for its robustness in classification tasks and low risk of overfitting. We implemented the model using the randomForest package [liaw2002] in R [R2021].
In the first step, we identified agro-climatic clusters by calculating dissimilarities at cocoa occurrence locations using 13 selected climate variables [shi2006]. We then grouped these locations using Ward hierarchical clustering, selecting the number of clusters based on visual inspection of the resulting dendrogram.
In the second step, we classified current climate conditions into the identified agro-climatic clusters. To reduce bias and exclude climatically unrealistic areas, we combined cocoa occurrence points with a background sample (1:1 ratio) drawn randomly from cocoa-producing countries [VanDerWal2009]. Each RF model was configured with 100 forests and 1000 trees, and repeated 25 times to ensure stability. We split the data into 80% for training and 20% for evaluation. The final classification maps represent the modal prediction across all trees, with “mixed” labels assigned to ambiguous locations and “limitations” for areas with partial suitability.
Results.We identified five distinct agro-climatic types across Ghana and Côte d’Ivoire, each showing unique patterns of cocoa suitability. To interpret and describe these types, we examined the confidence intervals of each climate variable across the RF-defined clusters (Supplementary Fig. S3). Based on their distinguishing climate features, we described the five agro-climatic types as follows:
  1. 1.
    Type 1: Moderate – very dry: Intermediate values for most temperature variables (Bio 3, 4, 8), but low values for precipitation during the driest month (Bio 14) and high number of consecutive months with less than 100 mm of precipitation (Bio 20).
  2. 2.
    Type 2: Hot & dry: Relatively high temperature values (Bio 8 and 11), high precipitation seasonality (Bio 15), long dry season (Bio 20).
  3. 3.
    Type 3: Hot & wet: Relatively high temperature values (Bio 8 and 11), high precipitation during the driest month (Bio 14), lowest number of consecutive dry months (Bio 20).
  4. 4.
    Type 4: Cooler & wet: Low mean temperature of the wettest quarter (Bio 8), high precipitation of the coldest quarter (Bio 19), relatively high number of consecutive dry months (Bio 20), high potential evapotranspiration (Bio 25).
  5. 5.
    Type 5: Very hot & very wet: High temperature seasonality and values (Bio 4, 8, and 11), low number of consecutive dry months (Bio 20).
These five agro-climatic types were mapped as five agro-climatic zones for Ivory Coast and Ghana, along with three additional classifications: "unsuitable" zones, which are unlikely to support cocoa due to climatic constraints, "mixed" zones which could not be clearly categorized into one of the agro-climatic zones and "limitations" zones which were classified unsuitable but were found to contain cocoa occurrences (Extended Data Figure Extended Data 1). This mapping provides a comprehensive view of the suitability of different areas for cocoa cultivation, considering the varying constellations of climatic attributes. Briefly, the Type 1 – moderate- very dry zone spanned the Northern cocoa zone from Yamoussoukro (CdI) to Ahafo (Ghana). Type 2 – hot & dry climate can be found South of Type 1 and North of Type 3 from Goh (CdI) to Ashanti (Ghana). The central cocoa zones Soubre and Divo in CdI, as well as Western Region in Ghana were characterized by Type 3 - hot & wet climate. Type 4 – cooler & wet – climate can be found in the Montagnes area (CdI) and the Ghanaian boarder with Togo, which also has some higher elevation area. Type 5 – very hot & very wet climate mostly covered coastal areas along the Gulf of Guinea (Extended Data Figure Extended Data 1).
Sampling design for shade-tree ground-truth data
To distribute our ground truth sampling effort in Ghana, we first randomly selected eight administrative districts within each of four agro-climatic zones. The four agro-climatic zones were defined earlier by Bunn et al. [bunn19], using a random forest machine learning algorithm to classify cocoa growing regions in Ghana according to climate and soil characteristics. Within each administrative district we randomly chose either one or two cocoa-growing communities, with the constraint that the selected communities have at least 25 recently mapped cocoa farms. We then randomly picked between 16 and 39 farms per community, ensuring that farms were distributed relatively evenly across the community, and that each individual farm was managed by a different farmer where possible.
Ground-truthing in Côte d’Ivoire was done in collaboration with a sampling program initiated by the Sustainable Cocoa Initiative Support Programme (SCISP) and the Sustainable Agicultural Supply Chains Initiative (SASI) of the German Development Agency (GIZ). These farms were distributed in the Sud Comoé and Indénié-Djuablin administrative regions in the east and Cavally region in the west of Côte d’Ivoire. Subsequent to this, we extended the agro-climatic zoning methods of Bunn et al.  [bunn19] to include Côte d’Ivoire (Supplementary Methods, Extended Data Fig. Extended Data 1), which indicated that these 60 farms were distributed across two of the four original cocoa agro-climatic zones identified in Ghana, and an additional cool and wet zone in the west of Côte d’Ivoire that was not originally identified in Ghana, where it only occurs in the Volta region (Extended Data Fig. Extended Data 1). We calculate all results with respect to the updated zones across both countries.
To identify the additional 69 farms with higher cover, we relied on farmer-reported shade-tree densities combined with interactive visual interpretation of Google Earth images. These additional high-cover farms were distributed in 30 cocoa-growing communities across four administrative districts in two agro-climatic zones.
Generating digital surface and terrain models
We used DroneDeploy software (DroneDeploy Inc., Version 2.236.0) to generate a digital surface model (DSM; top of canopy) and a digital terrain model (DTM) and subtracted the two to obtain a canopy height map for each farm. The DSM is generated first using the drone imagery, which captures all visible surface features, including buildings, vegetation, and other structures. The point cloud data from the drone imagery is also analyzed to detect and classify ground points, of which there tended to be many identified for each farm. To create the DTM, the algorithm filters out all non-ground elements, isolating ground points to represent the bare-earth surface. In areas where direct ground visibility is limited, interpolation techniques are applied to estimate terrain elevation based on surrounding data. The final result is a continuous surface, though some degree of approximation is involved depending on terrain complexity and data density. We note that individual farms are relatively flat, which favors more accurate terrain interpolation based on available ground points, and also limits the chance of systematic biases in terrain estimation that could systematically bias our canopy height estimates within or between farms. Finally, our threshold height for distinguishing cocoa from shade trees (see below) is conservative, such that our methods are likely to be robust to any (likely small in both magnitude and spatial extent) positive or negative mis-estimates of terrain elevation within farms.
Threshold height calculations
To determine a suitable height threshold to distinguish between shade trees and cocoa trees (which tend to be relatively short), we manually labelled 6800 polygons delimiting small sections of cocoa monocultures (total area 118 ha) in 230 cocoa farms distributed across the study area. These monoculture areas were identified using visual interpretation of orthophotos derived from the drone images. For all labelled polygons we then extracted the height of cocoa trees from the canopy height map. This analysis showed that 99.7% of all cocoa trees remained below a height of eight meters (Supplementary Fig. S4), while previous work shows that heights of shade trees tend to be higher than eight meters [blaser_2021]. We therefore used a threshold height of eight meters; vegetation above this height could be confidently attributed to shade trees rather than cocoa trees.
Shade-cover model performance
On our hold-out test set of 259 farms, we achieve a mean absolute error of 5.77 percentage points (pp.), a root mean squared error of 8.22 pp., and a mean (bias) error of 1.63 pp. We visualize a residual analysis for different reference shade cover values in Supplementary Fig. S1a).
Comparison to existing AGBD estimates
We compare our biomass map with the European Space Agency’s (ESA) Climate Change Initiative (CCI) state-of-the-art Biomass map for 2018 (ESA CCI map). The ESA CCI map is intended to provide global coverage of AGBD estimates with a ground sampling distance of $100\times 100$ meters (i.e., four times less pixels per area than our map), with the aim of monitoring global biomass change in the context of climate change. The ESA CCI map is created bi-annually based on multi-modal input data like spaceborne SAR (newer versions also include GEDI) and processed using a proprietary algorithm. Because we are lacking an independent, large-scale ground truth dataset of AGBD collected in the field across our region of interest, we compare both maps to 44,200 GEDI footprints that fall in our hold-out test areas (i.e., unseen by our model during training and validation) and contain cocoa according to Kalischek et al.[kalischek23natfood]. One should keep in mind though that our map has been specifically fine-tuned to our area of interest (as opposed to the ESA CCI map, which is a global product) and it is using GEDI footprint-level AGBD data for training, too. Unknown global biases that could be present in the GEDI AGBD footprint-level product would thus most likely impact our map but not show up in the comparison to the hold-out GEDI footprints here, if those would exhibit the same kind of possible bias. This comparison shall therefore not judge the quality of the ESA CCI map but rather provide some rough notion about our map’s properties. Overall, our map achieves (all values in tonnes ha-1) a mean absolute error (MAE) of 36.2 (vs. ESA’s 56.9), a root mean squared error (RMSE) of 63.2 (vs. ESA’s 87.9) and a mean error (bias) of 6.5 (vs. ESA’s 33.11). This shows that, in cocoa areas, our map does not only exhibit lower errors than the comparison map, but is also less biased towards lower AGBD values. In Supplementary Fig. S1b, we additionally show an analysis of the residuals between the two maps and the GEDI reference footprints, to better understand the biases of both maps over different AGBD intervals.
Bayesian regression methods for quantifying relationship between cover and biomass
We quantified the relationship between shade-tree cover and total biomass with Bayesian linear regression, using our pixel-level estimates of shade-tree cover as the predictor variable, and our pixel-level estimates of total aboveground biomass as the response variable. Only pixels in cocoa growing areas with shade levels up to 40% cover were used for the analysis. That set is representative for the range of plausible agroforestry scenarios given current recommendations, and includes 99.89% of all data points. In addition to shade-tree cover, we included a second-order polynomial term (i.e., the predictor "shade-tree-cover2") to account for potential curvilinearity in the relationship between cover and biomass. We used a Gaussian likelihood function for the response variable, and for each model coefficient we used improper flat priors (a uniform distribution over the real numbers), a common uninformative prior in Bayesian analyses. The model was fitted using Stan (https://mc-stan.org) together with the R package brms[bürkner2017]. It was trained on a 10% random sample of the filtered dataset and saved as an .rds object to support reproducibility and future analysis. The model was fitted using four Markov chain Monte Carlo (MCMC) chains, each with 2000 iterations, with a warm-up period of 1000 iterations. This gave a total of 4000 samples. Trace plots were visually assessed to ensure model convergence, and we also checked that the scale reduction factor ($\hat{R}$) was approximately equal to one. We note that we did attempt to fit an error-in-variables model to account for measurement uncertainty in the predictor, but this model consistently failed to converge, so we proceeded with the simpler Bayesian regression model as described above. With respect to predicting carbon-sequestration potential the simpler model should be more conservative as, all else being equal, it will tend to underestimate the strength of the relationship between cover and biomass.
Additional carbon-sequestration calculations